Presentation is loading. Please wait.

Presentation is loading. Please wait.

ANU Shutdown Status AEM meeting Cons Gattuso 6/3/2012.

Similar presentations

Presentation on theme: "ANU Shutdown Status AEM meeting Cons Gattuso 6/3/2012."— Presentation transcript:

1 ANU Shutdown Status AEM meeting Cons Gattuso 6/3/2012

2 NuMI Activities Target Pile dehumidifier repairs & upgrades - 95% complete –Electrical hook ups remain should be complete today or tomorrow Phase 1 of Horn 2 RAW system NoVA/ANU upgrade -100% complete New NoVA/ANU Target Pile fan cooling coil 75% complete –New cools are on order and will arrive in 6-8 weeks Un-forseen problem with existing cools start to leak after system was being refilled. Relocation of TP fan controls and sum box -100% complete –This move was require to open space for Raw water system upgrades Work cell lighting & electrical service installation -90% complete –What remain is some electrical installation for decay pipe De humidifier cooling coil. Facility drain & drip ceiling cleaning-100% complete (will need to be repeated at the end of shut down) –Complete drain cleaning in shaft area –Completed drain cleaning in the Absorber Passageway & Hall

3 4 NuMI Activities Replaced Pretarget magnets Q113, Q114, & Q120 90 % Complete –Vacuum has been restored –Cable pulls are complete Terminations and hook up will occur later this week Pull 6 electrical cables from the shaft to new quad locations -50% complete –Complete cable pulls through the upper labyrinth –Started cables pulls in the Pre-Target area –Installed handrail above emergency walkway –Installed scaffold in PT area Chase to be opened in week 8 (~ 2weeks from now)

4 Decommissioning MI/RR Stochastic Cooling system – 100 % Complete –101-103 Tanks and Stands removed –210-212 Tanks and Stands removed Spool piece in place Bake setup underway Bake Schedule for Tuesday afternoon after Service Building power outages R22 Line - 90 % Complete –Vacuum components in line have been harvested –Instrumentation components have been harvested BPM Toroids Multi Wires –Recycler Lambertson at 212 location Waiting on spool piece will be ready this week –MI Lambertson at 221 was removed (class 2 10-100 mr/hr)

5 Decommissioning MI/RR R32 Line - 25 % Complete –Vacuum components in line will need to be harvested –Instrumentation components will need to be harvested BPM –Toroid's –Multi Wires –Lambertson removal MI 321 location- 100 % Complete (was class 3 100-200 mr/hr) –Spool piece in its place RR 328 location – Schedule for later this week –Correction elements 100 % Complete Have been harvested Already reassigned to new girders –Mirror magnet MGS-001 Removed and relocated for future installation in the new extraction line

6 Decommissioning MI/RR August 31, 2010 DOE Review MI-620 Region A150 line Decommissioning - 100 % Complete –Two 2Q130 were removed from 902 and 901 location Magnets are being reworked for NuMI beam line spares Required the removal and reinstallation of Q618 –MI 620 Lamberstons Two have been removed and placed in short term storage (MSB) One Lambertson Staged for future installation at the 306 location –MI 620 wide aperture quad has been swopped out with 306 Quad –V901 C-Magnet was removed and put into viable spares pull (TPL) –2 kickers left in the A50 alcove (class 2-3) Replaced of NuMI stub 100 % Complete –Replaced Q111 & Q112 (3Q120 magnets Peter porch) –Reworked Stand for V109 to abate future install problems for mangets in the stub region

7 MI/RR RF section MI RF system Decommissioning - 100 % Complete (Shipped to TPL for storage) –Six 2.5 MHz RF cavities were removed –One 106 MHZ RF cavity were removed –Two Barrier Bucket RF cavities were removed MI RF system Installation/ Maintenance –RF4A cavity installation -100 % Complete –RF15A cavity installation -20% Complete Space for cavity has been made –Cavities determined to need tunnel repairs (1 to 2 weeks/ Cavity) We need to pull RF cavities 8 and 16 to replace bellows and flanges. Water cooling loop work for cavities 7, 12, 13, 15, and 17

8 Cables Pulls/Misc. Cable Pulls (~450 K feet of cables) –MI-50 House BPM cables 100 % Complete –5 days to complete – re-spooled to length –Historically ~2 week Sextupole cables 100 % Complete –MI-40 House BPM cables –Being staged –Pulls will start Tuesday BPM preamp- ion clearing module removal 100 % Complete –1 55 gallon barrel of circuit boards (mixed wasted) –1 64 cu. ft cube of rad waste August 31, 2010 DOE Review

9 Injection line Installation Recycler line installation -20% complete (Stand installation/relocation process halted) –Problems encountered Installation drawings slow to materialize Stands placement Interferences between the two beam lines Stands placement issues with respect to beam center line –Verify stake out marker locations –Verifying Lattice files –We have a plan……. MI line relocation –MI Quads and PDD magnets 60% complete MI-14 Fluorinert Piping Installation –Installed the brackets on the wall for the Fluorinert lines in the Q104 region –Relocated the Fluorinert manifold from the Q400 region to the Q104 region. –All the welds, from the penetration to the end of the Fluorinert run, that is, past the gate near the Q104 region, have been completed

10 Original Schedule

11 Current Schedule

Download ppt "ANU Shutdown Status AEM meeting Cons Gattuso 6/3/2012."

Similar presentations

Ads by Google